a plane moves from a town X to a town Y 30 km apart in the direction 60° and then changes direction moving to a point Z in the direction 120° if the point Z, is due east of X, find the distance of the point Z from point X.
1
Expert's answer
2019-09-16T09:48:44-0400

In order to have the complete problem description, whenever a direction of flight (e.g. in degrees) is given, it should also be mentioned in reference to which axis/bearing the direction (i.e. angle) is.


If the aircraft reroutes from "60\degree" to "120\degree" and eventually arrives to a town due east from the origin, we can deduct that it is initially "60\degree" "east of south", whereas the latter angle "120\degree" "east of south" would in other words mean "(180-120)\degree = 60\degree" "east of north".


With the first flight X to Y the displacement along "east" axis, in the east direction,

"\Delta s_{XY} = 30\text{ km }\cdot\sin(60\degree)"


Since then the aircraft flies at the same degree ("60\degree") with respect to "east of north", it compensates for the same opposite displacement along the "north-south" axis, thus arriving at a town Z due east of X, therefore displacement Y to Z along the "east" axis is symetrically the same as from X to Y:

"\Delta s_{YZ} = \Delta s_{XY} = 30\text{ km }\cdot\sin(60\degree)"


Result:

"\Delta s_{XZ} = \Delta s_{XY} + \Delta s_{YZ} = 2\Delta s_{XY} = 2\cdot30\text{ km }\cdot\sin(60\degree) = 30\sqrt{3}\text{ km } \approx 52\text{ km}."